Q 1 a) The Output and the total cost data for a firm are given below. Work out the following costs.


Total Fixed cost, Total Variable cost, Average Fixed cost, Average Variable cost, Average Total cost and


Marginal cost at various levels output.


Units of output 0 1 2 3 4 5 6


Total cost (Rs.) 60 90 100 105 115 135 180


b) If the total cost is positive at zero level of output, what does it signify?

1
Expert's answer
2022-02-16T12:48:29-0500

A)"FC=TC-VC" But when output is zero;

"VC=0" , since fixed costs does not change throughout the production, its equal to,

FC=60 60 60 60 60 60 60


"VC=TC-FC"

= 0 30 40 45 55 75 120

"AVC= \frac{VARIABLE COST}{TOTAL PRODUCT}"

=0 5 6.6667 7.5 9.1667 15 25


"AFC=\frac{FIXED COST}{TOTAL PRODUCT}"


=19 10 10 10 10 10 10

"ATC=AFC+AVC"


=10 15 16.6667 17.5 19.1667 25 30


"MC=\frac{CHANGE IN T0TAL COST}{CHANGE IN TOTAL PRODUCT}"

=- 30 10 5 10 20 45


B) This signify that there are no variable costs incurred but there are fixed costs incurred.